OpenStudy (anonymous):

Mr. and Mrs. Snow had four children in two year intervals. The sum of the ages of the children is 36. List the ages of the children from youngest to oldest.

OpenStudy (anonymous):

36=a+b+c+d , b=2+a , c=b+2 , d=c+2

OpenStudy (anonymous):

if you solve for the system you will get a = 6, b = 8, c = 10, d = 12

OpenStudy (anonymous):

Jerry has a total of 60 quarters and pennies in his pocket. These coins total $3.00. How many quarters are in Jerry's pocket?Jerry has a total of 60 quarters and pennies in his pocket. These coins total $3.00. How many quarters are in Jerry's pocket?

OpenStudy (anonymous):

3=.25x+.01y 60=x+y I wanna say x = 10., y = 50 10 quarters 50 pennies